If | -2X + 9 | = 3 then the possible value of
| -X | - X^2 would be :
A)30 B)-30. C)-42. D)42

Answer by rothauserc(4718)

| -2X + 9 | = 3
we have two cases to evaluate
:
1) -2X +9 = 3
-2X = -6
X = 3
:
2) -(-2x+9) = 3
2X -9 = 3
2X = 12
X = 6
:
now consider | -X | - X^2
:
1) For X=3
3 - 3^2 = -6
:
2) For X=6
6 - 36 = -30
:
******************************
The answer is B